OpenStudy (anonymous):

A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t^2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

@rebeccaskell94

OpenStudy (asnaseer):

when it hits the ground, what will h equal?

OpenStudy (anonymous):

0?

OpenStudy (asnaseer):

correct, so you need to solve:\[0=rt-16t^2\]

OpenStudy (asnaseer):

and you are given r=48

OpenStudy (anonymous):

so 0 = 48t - 16t^2?

OpenStudy (asnaseer):

correct - now factorise this and the answer should become clear

OpenStudy (anonymous):

how do i factorize this one i've always been bad at factoring :l

OpenStudy (asnaseer):

ok - look at each part. first lets concentrate on just the numbers 48 and 16 - what is common between 48 and 16?

OpenStudy (asnaseer):

what divides into both 48 and 16

OpenStudy (anonymous):

2, 4, 8, and 16

OpenStudy (asnaseer):

good - so the largest number that divides into both is 16

OpenStudy (anonymous):

yea

OpenStudy (asnaseer):

so we can take out 16 to get:\[0=16(3t-t^2)\]

OpenStudy (asnaseer):

now what goes into both t and \(t^2\)?

OpenStudy (anonymous):

t?

OpenStudy (asnaseer):

correct again! so we can take out a t to get:\[0=16t(3-t)\]

OpenStudy (asnaseer):

next, whenever something times something else equals zero, then you know that one of them must be equal to zero - agreed?

OpenStudy (anonymous):

yea

OpenStudy (asnaseer):

so we have:\[0=16t(3-t)=16t\times(3-t)\]therefore we can say that either:\[16t=0\]or:\[3-t=0\]

OpenStudy (anonymous):

alright

OpenStudy (asnaseer):

we therefore have two solutions for t - can you work them out?

OpenStudy (anonymous):

so t = 3 and t = 0?

OpenStudy (asnaseer):

well done! so we know the ball was on the ground when t=0 (i.e. at the start) and when t=3 (i.e. 3s after he kicked it in the air

OpenStudy (asnaseer):

does it make sense?

OpenStudy (anonymous):

yep

OpenStudy (asnaseer):

ok - good

OpenStudy (anonymous):

wait 1 more thing

OpenStudy (asnaseer):

yeah?

OpenStudy (anonymous):

the answer is one of those fill ins so what should i write? t = 3?

OpenStudy (asnaseer):

in the question it says: After how many seconds does the ball hit the ground? so the answer must be t=3

OpenStudy (asnaseer):

t=0 is just the starting point

OpenStudy (anonymous):

alright thanks

OpenStudy (asnaseer):

yw